OpenStudy (anonymous):

Find the Taylor Series of the function f(x)= (1/x) c=4 ..... and the interval of convergance.. neeed help please ...

OpenStudy (anonymous):

\[1/x = \sum_{n=0}^{\infty}(-1)^n (x-1)^n\] for \[\left| 1-x \right| <1\] Expansion at c=4: \[1/4-(x-4)/16+(x-4)^2/64-(x-4)^3/256+...\]